OpenStudy (anonymous):

After a computer is discounted by 20%, its price is $960. What was the original price of the computer before the discount?

OpenStudy (anonymous):

80% = $960 then 100% = 100/80*960 = $1,200

OpenStudy (anonymous):

960/(1-0.2) = 1200

OpenStudy (anonymous):

960(1-20%)=1200